OpenStudy (anonymous):

A company sells cardboard that is eight and seven-eighths millimeters thick. Write an equation to solve for the number of sheets of cardboard, c, in a pile that is thick. What must be done to both sides to solve the equation?

OpenStudy (yolo_king):

JW.org ik that website

OpenStudy (anonymous):

Thats awesome but can u please answer my question

OpenStudy (anonymous):

Me too hi !!!!

OpenStudy (yolo_king):

8 and 7/8 mm = (71/8) mm

OpenStudy (yolo_king):

372 and 3/4 mm = (1491/4) mm

OpenStudy (yolo_king):

(71/8) * c = (1491/4)

OpenStudy (yolo_king):

multiply both sides by (8/71)